# MIGHTY: Hermite Spline-based Efficient Trajectory Planning

### v3

@@ I Introduction
MIGHTY (proposed) Hermite spline Hermite cntrl pts, time. joint full polynomial+time space yes direct & local
@@ I-A Review of Unconstrained Frameworks
Decoupling into two stages relaxes hard constraints for unconstrained formulations. Richter et al. [17] use 9th-order polynomials with endpoint derivatives: a QP minimizes snap, then gradient-based time allocation adjusts duration. Endpoint derivatives are decision variables, so there is direct derivative parameterization. However, the free derivatives are globally coupled, so there is no local control. Collision avoidance inserts mid-segment waypoints, triggering repeated solves. Extensions [1, 14] improve stability and enable real-time replanning via ESDF-based potential costs.
Another standard unconstrained framework uses B-splines for trajectory representation, with time-related variables encoded in a knot vector. B-splines offer inherent smoothness and geometric local control: changing one control point affects only a bounded neighborhood of the curve. The convex hull property enables efficient collision checking and enforces dynamic feasibility [18, 23, 25, 24]. However, B-splines do not provide direct local control of higher-order derivatives. Derivatives at any curve point are implicit functions of multiple neighboring control points. One can bound derivatives efficiently via the convex hull property, but cannot independently prescribe a specific velocity or acceleration at a given knot through a single variable. This geometric local control property has been widely exploited for efficient trajectory refinement and replanning in cluttered environments [6, 15, 20, 4]. Zhou et al. [23] parameterize trajectories as non-uniform cubic B-splines initialized from kinodynamic A* paths. The nonlinear optimization minimizes the cost considering smoothness, a repulsive collision potential from an Euclidean distance field (EDF), and soft penalties on derivative control points exceeding maximum velocity and acceleration. Finally, an iterative time-adjustment method rescales knot spans of the non-uniform spline to guarantee dynamic feasibility. EGO-Planner [25] and RAPTOR [24] simplify the representation by using uniform B-splines with fixed knot spans, leveraging convex hull properties to enforce velocity, acceleration, and jerk constraints directly on control points. EGO-Planner optimizes trajectories via a quasi-Newton solver (L-BFGS) [9] for fast gradient-based convergence, with an additional time-reallocation and anisotropic curve-fitting refinement stage that resizes knot spans and fits a new spline to ensure feasibility and smoothness. RAPTOR also employs a two-stage process: a closed-form QP warm-start for initial smoothness, followed by a nonlinear refinement phase that penalizes collision via an ESDF and ensures dynamic limits via control point constraints.
@@ I-B MIGHTY Contributions
    Hermite Spline Planner: Joint spatiotemporal optimization in a single solve, with explicit parameterization of positions and derivatives at each knot. This representation enables local control of higher-order dynamics and trajectory shape, which the optimizer can leverage when needed. The approach, MIGHTY, achieves fast solve times and superior performance.
@@ II Hermite Spline
This section introduces the Hermite-spline notation used in MIGHTY’s formulation. MIGHTY uses a degree-$d$ Hermite spline with $M$ segments. For odd degree $d=2k+1$, each segment is parameterized by the knot values of the position and its derivatives up to order $k$. For example, a quintic Hermite spline ($d=5$, $k=2$) specifies position, velocity, and acceleration. Across knots, a Hermite spline automatically guarantees $\mathcal{C}^{k}$ continuity by sharing the knot values. Although Hermite splines apply in any dimension, we work in $\mathbb{R}^{3}$ for quadrotor planning and use $d=5$ in this paper. For each segment $s=0,\dots,M-1$, we define a normalized time variable $\tau_{s}\in[0,1]$ on segment $s$ as $\tau_{s}=(t-t_{s})/T_{s}\,\in[0,1].$ Then each segment can be denoted as: $\mathbf{x}(t)=\mathbf{x}_{s}(\tau_{s}),\,t\in\bigl[t_{s},\;t_{s+1}\bigr],$ where $t_{s}:=\sum_{r=0}^{s-1}T_{r}$, and $T_{s}$ is the duration of segment $s$. Then we write
@@
where $\mathbf{H}_{s,0}=\mathbf{p}_{s},\mathbf{H}_{s,1}=T_{s}\,\mathbf{v}_{s},\mathbf{H}_{s,2}=\tfrac{1}{2}T_{s}^{2}\,\mathbf{a}_{s},\mathbf{H}_{s,3}=\mathbf{p}_{s+1},\ \mathbf{H}_{s,4}=T_{s}\,\mathbf{v}_{s+1},\ \mathbf{H}_{s,5}=\tfrac{1}{2}T_{s}^{2}\,\mathbf{a}_{s+1}.$ where $h_{k}(\tau_{s})$ are the standard quintic Hermite basis functions. The $\mathbf{p}_{i},\mathbf{v}_{i},\mathbf{a}_{i}\in\mathbb{R}^{3}\;(i=0,\dots,M)$ and $T_{s}>0\;(s=0,\dots,M-1)$ are the interior knot positions, velocities, accelerations, and segment durations, respectively.
@@ III-A Decision Variables
As shown in Section II, the knot states $\mathbf{p}_{i},\mathbf{v}_{i},\mathbf{a}_{i}(i=0,\dots,M)$ and segment durations $T_{s}(s=0,\dots,M-1)$ fully determine the trajectory. Thus we optimize the interior positions, velocities, accelerations, and per-segment durations:
$\mathbf{z}=\bigl[\,\mathbf{p}_{1}^{\top},\mathbf{v}_{1}^{\top},\mathbf{a}_{1}^{\top},\ldots,\mathbf{p}_{M-1}^{\top},\mathbf{v}_{M-1}^{\top},\mathbf{a}_{M-1}^{\top},T_{0},\ldots,T_{M-1}\bigr].^{\top}$
@@
To keep the optimization unconstrained while enforcing certain requirements (e.g., $T_{s}>0$), we apply diffeomorphism, such as $T_{s}\;=\;\phi(\sigma_{s})$ with $\phi(\sigma)=e^{\sigma},\ \sigma_{s}\in\mathbb{R}.$ We also find that optimizing raw derivative knots $(\mathbf{v}_{i},\mathbf{a}_{i})$ can be numerically unstable. Section III-D introduces scalings that improve stability without changing the optimum. For clarity, the formulas below are written in terms of the original variables $(\mathbf{p}_{s},\mathbf{v}_{s},\mathbf{a}_{s},T_{s})$; applying a diffeomorphism only inserts chain-rule factors in the gradients.
@@ III-C Objective and Closed-Form Gradient
To make the formulation basis-independent, we express the objective and its gradients w.r.t a generic control-point (or coefficient) vector $\mathbf{c}$, independent of the chosen parameterization (Hermite end states, Bézier control points, etc.). As a concrete example, the affine map between Hermite and Bézier on segment $s$ is
@@
And $\partial\mathbf{c}_{s}/\partial T_{s}$ ($\nabla_{T_{s}}\mathbf{c}_{s}$) used in the coefficient path in Eq. (4) can be computed in closed form as
@@
Therefore, to obtain the gradient, we need to compute $\mathbf{g}_{\mathbf{c}_{s}}$ and $\bigl.\partial J_{s}/\partial T_{s}\bigr|_{\text{explicit}}.$ Here we present two types of cost terms: (1) cost terms defined on control points, and (2) cost terms defined via sampled states along the trajectory, and their gradients used in Section IV.
@@ III-C1 Cost on Control Points
where $C_{s}=3600\,T_{s}^{-5}$. Define the (scalar) Bernstein-quadratic Gram matrix $G\in\mathbb{R}^{3\times 3}$ by $G_{mn}\;\triangleq\;\int_{0}^{1}B_{m}^{2}(\tau_{s})\,B_{n}^{2}(\tau_{s})\,d\tau_{s}\quad(m,n=0,1,2),$ so that, with $\Delta_{s}\triangleq$ then $\Delta_{s,0}^{\top},\Delta_{s,1}^{\top},\Delta_{s,2}^{\top}]^{\top}\in\mathbb{R}^{9}$,
@@ III-D Reparameterizations for Derivative Variables
and set $\hat{\mathbf{v}}_{i}=\bar{T}_{i}\,\mathbf{v}_{i},\quad\hat{\mathbf{a}}_{i}=\bar{T}_{i}^{2}\,\mathbf{a}_{i}.$ This reparameterization aims to reduce the explicit $1/T_{s}$ and $1/T_{s}^{2}$ factors’ effect in the gradients w.r.t. the derivative decision variables by optimizing the scaled variables $\{\hat{\mathbf{v}}_{i},\hat{\mathbf{a}}_{i}\}$ instead of $\{\mathbf{v}_{i},\mathbf{a}_{i}\}$. At the endpoints, the scaling uses the adjacent segment duration (no averaging), while at interior knots we use the local average duration to balance the influence of the two neighboring segments; in that case, the inverse-duration effects are mitigated but cannot be canceled for both segments simultaneously unless their durations match. Our ablation study in Section IV-E confirms the expected improvement in performance.
@@ IV-A Representation Benchmarking: Simple Case
We first benchmark MIGHTY against GCOPTER (which uses MINCO [22]) on a simple corner-avoidance task with fixed start and goal (Fig. 1). For a fair comparison, we implemented GCOPTER’s optimization cost in MIGHTY and use the same optimizer (L-BFGS [9]) and stopping tolerances. Both methods start from the same initial guess.
GCOPTER represents each segment as a fifth-order polynomial, so MINCO minimizes jerk (third-order effort) by construction. MIGHTY uses a fifth-order Hermite spline to match the polynomial degree. GCOPTER applies diffeomorphic parameterizations for waypoints and durations (waypoints remain inside the overlap of safe-flight corridors; durations stay positive). The only difference is that MIGHTY includes an explicit jerk smoothness term (Sec. III-C1), while GCOPTER does not add a separate smoothness cost because MINCO already minimizes jerk.
Figure 1: Benchmarking environment for a simple corner avoidance scenario. The blue-shaded polygons show the safe flight corridor. MIGHTY’s trajectory is colored by speed (warmer colors indicate faster speeds).
Figure 2: Benchmarking results for MIGHTY and GCOPTER for simple corner avoidance scenario in Fig 1. MIGHTY achieves lower computation time, travel time, and path length, while GCOPTER yields lower jerk.
@@
The result in Fig. 2 shows that MIGHTY achieves a shorter travel time, with slightly faster computation and shorter path than GCOPTER. While GCOPTER produces smoother trajectories with lower jerk due to its constraint within the MINCO class, MIGHTY explores a larger search space to identify faster and more effective trajectories. This flexibility enables MIGHTY to achieve superior performance, and Section IV-B further investigates the jerk-performance trade-off.
@@ IV-D Representation Benchmarking: Complex Case
Figure 5: Complex-scene benchmarking setup. MIGHTY’s trajectory is color-mapped by speed (warm=fast), and GCOPTER’s trajectory is blue. The blue-shaded polygons show the safe flight corridor. The start is at $(0,0,0.5)\,$\mathrm{m}$\$, and 24 goals lie on a grid with $x,y\in[-15,15]\,$\mathrm{m}$\$ and $z=$2.5\text{\,}\mathrm{m}$\$. Blue-shaded polygons show the safe flight corridor sequence from one run.
@@ IV-E Ablation Study on Scaled and Unscaled Variables
As discussed in Section III-D, we optimize over scaled variables to improve numerical stability. We evaluate this via ablation using the same complex benchmarking setup. Fig. 6 shows the scaled variant is $\approx 2\times$ faster with lower jerk, while path length and travel time are nearly identical, confirming that scaling improves efficiency without degrading quality.
@@ IV-F Benchmarking in Static Environments
$R_{\mathrm{succ}}$ [%] $T_{\mathrm{opt}}$ [ms] $T_{\mathrm{total}}$ [s] $T_{\mathrm{trav}}$ [s] $L_{\mathrm{path}}$ [m] $S_{\mathrm{jerk}}$ [m/s2] $\bar{S}_{\mathrm{jerk}}$ [m/s3] $\rho_{\mathrm{vel}}$ [%] $\rho_{\mathrm{acc}}$ [%] $\rho_{\mathrm{jerk}}$ [%]
@@
To thoroughly evaluate MIGHTY, we benchmark it in an obstacle-rich environment (Fig. 7) against EGO-Swarm [27], EGO-Swarm2 [26], and SUPER [16]. Static cylindrical obstacles are placed randomly with radii sampled in $1.0\text{\,}\mathrm{m}$,$1.5\text{\,}\mathrm{m}$\$\,] and height $6.0\text{\,}\mathrm{m}$, occupying a $300\text{\,}\mathrm{m}$\times$40\text{\,}\mathrm{m}$$ area. The agent starts at $(0,0,3)\,$\mathrm{m}$\$ and the goal is $(305,0,3)\,$\mathrm{m}$\$. We constrain velocity, acceleration, and jerk to $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, $a_{\max}=$10.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, and $j_{\max}=$30.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{3}$\$. For global planning we run A*, then build a safe-flight corridor (SFC) [10] for local optimization. The intermediate waypoints for A* are projected from the terminal goal onto the local occupancy map centered on the drone. The average computation time for A* and SFC generation is $0.26\text{\,}\mathrm{m}\mathrm{s}$ and $0.48\text{\,}\mathrm{m}\mathrm{s}$, respectively.
For soft constraints we use a smooth hinge $\phi_{\mu}(\cdot)$. Time integrals are evaluated by trapezoidal quadrature with $\kappa$ samples per segment unless a closed form is available. Safe-flight corridors (SFCs) are modeled per segment $s$ as the intersection of halfspaces
$\text{SFC}_{s}\;\triangleq\;\big\{\mathbf{x}\in\mathbb{R}^{3}\;\big|\;\mathbf{a}_{s,h}^{\top}\mathbf{x}\leq b_{s,h}-C_{\text{SFC}},\ \ h=1,\dots,H_{s}\big\},$
@@
where $J_{\text{SFC}}\;=\;\sum_{s}\int_{0}^{T_{s}}\sum_{h=1}^{H_{s}}\phi_{\mu}\!\big(\,\mathbf{a}_{s,h}^{\top}\mathbf{x}(t)-b_{s,h}+C_{\text{SFC}}\,\big)\,dt,$ penalizes leaving the SFC, $J_{v}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{v}(t)\|^{2}-v_{\max}^{2})\,dt$, $J_{a}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{a}(t)\|^{2}-a_{\max}^{2})\,dt$, and $J_{j}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{j}(t)\|^{2}-j_{\max}^{2})\,dt$, Note that the jerk smoothness cost $\tilde{J}_{\text{smooth}}$ uses the closed form given in Eq. 6, which avoids per-sample recomputation. MIGHTY’s weights are $w_{T}=$5\text{\times}{10}^{2}$\$, $w_{\mathrm{smooth}}=${10}^{-1}$\$, $w_{\text{SFC}}=${10}^{3}$\$, $w_{v}=${10}^{3}$\$, $w_{a}=${10}^{3}$\$, $w_{j}=${10}^{3}$\$. We run 10 trials per method. Metrics: $R_{\mathrm{succ}}$ [%] (success rate; collision-free, reaches goal); $T_{\mathrm{opt}}$ [ms] (local optimization time); $T_{\mathrm{total}}$ [ms] (total planning time); $T_{\mathrm{trav}}$ [s] (travel time); $L_{\mathrm{path}}$ [m] (path length); $S_{\mathrm{jerk}}=\!\int\!\lVert\mathbf{j}(t)\rVert\,dt$ [m/s3] (smoothness); $\bar{S}_{\mathrm{jerk}}=\sqrt{\tfrac{1}{T}\int\!\lVert\mathbf{j}(t)\rVert^{2}\,dt}$ [m/s3] (RMS jerk; time-normalized); $\rho_{\mathrm{vel}},\,\rho_{\mathrm{acc}},\,\rho_{\mathrm{jerk}}$ [%] (velocity/acceleration/jerk violations). We report $\bar{S}_{\mathrm{jerk}}$ to account for differing travel times across methods.

### v4

@@ I Introduction
MIGHTY (proposed) Hermite spline Hermite cntrl pts, time joint full polynomial+time space yes direct & local
@@ I-A Review of Unconstrained Frameworks
Decoupling into two stages enables unconstrained formulations by replacing hard constraints with soft penalties. Richter et al. [11] use 9th-order polynomials with endpoint derivatives: a QP minimizes snap, then gradient-based time allocation adjusts duration. Endpoint derivatives are decision variables, so there is direct derivative parameterization. However, the free derivatives are globally coupled, so there is no local control. Collision avoidance inserts mid-segment waypoints, triggering repeated solves. Extensions [16, 17] improve stability and enable real-time replanning via ESDF-based potential costs.
Another standard unconstrained framework uses B-splines for trajectory representation, with time-related variables encoded in a knot vector. B-splines offer inherent smoothness and geometric local control: changing one control point affects only a bounded neighborhood of the curve. The convex hull property enables efficient collision checking and enforces dynamic feasibility [4, 6, 7, 9]. However, B-splines do not provide direct local control of higher-order derivatives. Derivatives at any curve point are implicit functions of multiple neighboring control points. One can bound derivatives efficiently via the convex hull property, but cannot independently prescribe a specific velocity or acceleration at a given knot through a single variable. This geometric local control property has been widely exploited for efficient trajectory refinement and replanning in cluttered environments [18, 19, 20, 21]. Zhou et al. [6] parameterize trajectories as non-uniform cubic B-splines initialized from kinodynamic A* paths. The nonlinear optimization minimizes the cost considering smoothness, a repulsive collision potential from a Euclidean distance field (EDF), and soft penalties on derivative control points exceeding maximum velocity and acceleration. Finally, an iterative time-adjustment method rescales knot spans of the non-uniform spline to guarantee dynamic feasibility. EGO-Planner [7] and RAPTOR [9] simplify the representation by using uniform B-splines with fixed knot spans, leveraging convex hull properties to enforce velocity, acceleration, and jerk constraints directly on control points. EGO-Planner optimizes trajectories via a quasi-Newton solver (L-BFGS) [22] for fast gradient-based convergence, with an additional time-reallocation and anisotropic curve-fitting refinement stage that resizes knot spans and fits a new spline to ensure feasibility and smoothness. RAPTOR also employs a two-stage process: a closed-form QP warm-start for initial smoothness, followed by a nonlinear refinement phase that penalizes collision via an ESDF and ensures dynamic limits via control point constraints.
@@ I-B MIGHTY Contributions
    Hermite Spline Planner: Joint spatiotemporal optimization in a single solve, with explicit parameterization of positions and derivatives at each knot. This representation enables local control of higher-order dynamics and trajectory shape, which the optimizer can leverage when needed. The approach, MIGHTY, achieves fast solve times and lower travel time than baselines.
@@ II Hermite Spline
This section introduces the Hermite-spline notation used in MIGHTY’s formulation. MIGHTY uses a degree-$d$ Hermite spline with $M$ segments. For odd degree $d=2\nu+1$, each segment is parameterized by the knot values of the position and its derivatives up to order $\nu$. For example, a quintic Hermite spline ($d=5$, $\nu=2$) specifies position, velocity, and acceleration. Across knots, a Hermite spline automatically guarantees $\mathcal{C}^{\nu}$ continuity by sharing the knot values. Although Hermite splines apply in any dimension, we work in $\mathbb{R}^{3}$ for quadrotor planning and use $d=5$ in this paper. For each segment $s=0,\dots,M-1$, we define a normalized time variable $\tau_{s}\in[0,1]$ on segment $s$ as $\tau_{s}=(t-t_{s})/T_{s}\,\in[0,1].$ Then each segment can be denoted as: $\mathbf{x}(t)=\mathbf{x}_{s}(\tau_{s}),\,t\in\bigl[t_{s},\;t_{s+1}\bigr],$ where $t_{s}:=\sum_{r=0}^{s-1}T_{r}$, and $T_{s}$ is the duration of segment $s$. Then we write
@@
where $\mathbf{H}_{s,0}=\mathbf{p}_{s},\mathbf{H}_{s,1}=T_{s}\,\mathbf{v}_{s},\mathbf{H}_{s,2}=\tfrac{1}{2}T_{s}^{2}\,\mathbf{a}_{s},\mathbf{H}_{s,3}=\mathbf{p}_{s+1},\ \mathbf{H}_{s,4}=T_{s}\,\mathbf{v}_{s+1},\ \mathbf{H}_{s,5}=\tfrac{1}{2}T_{s}^{2}\,\mathbf{a}_{s+1}.$ Here, $h_{k}(\tau_{s})$ are the standard quintic Hermite basis functions. The $\mathbf{p}_{i},\mathbf{v}_{i},\mathbf{a}_{i}\in\mathbb{R}^{3}\;(i=0,\dots,M)$ and $T_{s}>0\;(s=0,\dots,M-1)$ are the interior knot positions, velocities, accelerations, and segment durations, respectively.
@@ III-A Decision Variables
As shown in Sec. II, the knot states $\mathbf{p}_{i},\mathbf{v}_{i},\mathbf{a}_{i}(i=0,\dots,M)$ and segment durations $T_{s}(s=0,\dots,M-1)$ fully determine the trajectory. Thus we optimize the interior positions, velocities, accelerations, and per-segment durations:
$\mathbf{z}=\bigl[\,\mathbf{p}_{1}^{\top},\mathbf{v}_{1}^{\top},\mathbf{a}_{1}^{\top},\ldots,\mathbf{p}_{M-1}^{\top},\mathbf{v}_{M-1}^{\top},\mathbf{a}_{M-1}^{\top},T_{0},\ldots,T_{M-1}\bigr]^{\top}.$
@@
To keep the optimization unconstrained while enforcing certain requirements (e.g., $T_{s}>0$), we apply diffeomorphism, such as $T_{s}\;=\;\phi(\sigma_{s})$ with $\phi(\sigma)=e^{\sigma},\ \sigma_{s}\in\mathbb{R}.$ We also find that optimizing raw derivative knots $(\mathbf{v}_{i},\mathbf{a}_{i})$ can be numerically unstable. Sec. III-D introduces scalings that improve stability without changing the optimum. For clarity, the formulas below are written in terms of the original variables $(\mathbf{p}_{s},\mathbf{v}_{s},\mathbf{a}_{s},T_{s})$; applying a diffeomorphism only inserts chain-rule factors in the gradients.
@@ III-C Objective and Closed-Form Gradient
To make the formulation basis-independent, we express the objective and its gradients w.r.t. a generic control-point (or coefficient) vector $\mathbf{c}$, independent of the chosen parameterization (Hermite end states, Bézier control points, etc.). As a concrete example, the affine map between Hermite and Bézier on segment $s$ is
@@
The coefficient gradient $\partial\mathbf{c}_{s}/\partial T_{s}$ ($\nabla_{T_{s}}\mathbf{c}_{s}$) used in the coefficient path in Eq. (4) can be computed in closed form as
@@
Therefore, to obtain the gradient, we need to compute $\mathbf{g}_{\mathbf{c}_{s}}$ and $\bigl.\partial J_{s}/\partial T_{s}\bigr|_{\text{explicit}}.$ Here we present two types of cost terms: (1) cost terms defined on control points, and (2) cost terms defined via sampled states along the trajectory, and their gradients used in Sec. IV.
@@ III-C1 Cost on Control Points
where $C_{s}=3600\,T_{s}^{-5}$. Define the (scalar) Bernstein-quadratic Gram matrix $G\in\mathbb{R}^{3\times 3}$ by $G_{mn}\;\triangleq\;\int_{0}^{1}B_{m}^{2}(\tau_{s})\,B_{n}^{2}(\tau_{s})\,d\tau_{s}\quad(m,n=0,1,2),$ so that, with $\Delta_{s}\triangleq[\Delta_{s,0}^{\top},\Delta_{s,1}^{\top},\Delta_{s,2}^{\top}]^{\top}\in\mathbb{R}^{9}$,
@@ III-D Reparameterizations for Derivative Variables
and set $\hat{\mathbf{v}}_{i}=\bar{T}_{i}\,\mathbf{v}_{i},\quad\hat{\mathbf{a}}_{i}=\bar{T}_{i}^{2}\,\mathbf{a}_{i}.$ This reparameterization aims to reduce the explicit $1/T_{s}$ and $1/T_{s}^{2}$ factors’ effect in the gradients w.r.t. the derivative decision variables by optimizing the scaled variables $\{\hat{\mathbf{v}}_{i},\hat{\mathbf{a}}_{i}\}$ instead of $\{\mathbf{v}_{i},\mathbf{a}_{i}\}$. At the endpoints, the scaling uses the adjacent segment duration (no averaging), while at interior knots we use the local average duration to balance the influence of the two neighboring segments; in that case, the inverse-duration effects are mitigated but cannot be canceled for both segments simultaneously unless their durations match. Our ablation study in Sec. IV-E confirms the expected improvement in performance.
@@ IV-A Representation Benchmarking: Simple Case
We first benchmark MIGHTY against GCOPTER (which uses MINCO [13]) on a simple corner-avoidance task with fixed start and goal (Fig. 1). For a fair comparison, we implement GCOPTER’s optimization cost in MIGHTY and use the same optimizer (L-BFGS [22]) and stopping tolerances. Both methods start from the same initial guess.
GCOPTER represents each segment as a fifth-order polynomial, so MINCO minimizes jerk (third-order effort) by construction. MIGHTY uses a fifth-order Hermite spline to match the polynomial degree. GCOPTER applies diffeomorphic parameterizations for waypoints and durations (waypoints remain inside the overlap of safe flight corridors (SFCs); durations stay positive). The only difference is that MIGHTY includes an explicit jerk smoothness term (Sec. III-C1), while GCOPTER does not add a separate smoothness cost because MINCO already minimizes jerk.
Figure 1: Benchmarking environment for a simple corner avoidance scenario. The blue-shaded polygons show the SFC. MIGHTY’s trajectory is colored by speed (warmer colors indicate faster speeds).
Figure 2: Benchmarking results for MIGHTY and GCOPTER for simple corner avoidance scenario in Fig. 1. MIGHTY achieves lower computation time, travel time, and path length, while GCOPTER yields lower jerk.
@@
The result in Fig. 2 shows that MIGHTY achieves a shorter travel time, with slightly faster computation and shorter path than GCOPTER. While GCOPTER produces smoother trajectories with lower jerk due to its constraint within the MINCO class, MIGHTY explores a larger search space to identify faster and more effective trajectories. This flexibility enables MIGHTY to achieve superior performance, and Sec. IV-B further investigates the jerk-performance trade-off.
@@ IV-D Representation Benchmarking: Complex Case
Figure 5: Complex-scene benchmarking setup. MIGHTY’s trajectory is color-mapped by speed (warm=fast), and GCOPTER’s trajectory is blue. The blue-shaded polygons show the SFC. The start is at $(0,0,0.5)\,$\mathrm{m}$\$, and 24 goals lie on a grid with $x,y\in[-15,15]\,$\mathrm{m}$\$ and $z=$2.5\text{\,}\mathrm{m}$\$.
@@ IV-E Ablation Study on Scaled and Unscaled Variables
As discussed in Sec. III-D, we optimize over scaled variables to improve numerical stability. We evaluate this via ablation using the same complex benchmarking setup. Fig. 6 shows the scaled variant is $\approx 2\times$ faster with lower jerk, while path length and travel time are nearly identical, confirming that scaling improves efficiency without degrading quality.
@@ IV-F Benchmarking in Static Environments
$R_{\mathrm{succ}}$ [%] $T_{\mathrm{opt}}$ [ms] $T_{\mathrm{total}}$ [ms] $T_{\mathrm{trav}}$ [s] $L_{\mathrm{path}}$ [m] $S_{\mathrm{jerk}}$ [m/s2] $\bar{S}_{\mathrm{jerk}}$ [m/s3] $\rho_{\mathrm{vel}}$ [%] $\rho_{\mathrm{acc}}$ [%] $\rho_{\mathrm{jerk}}$ [%]
@@
To thoroughly evaluate MIGHTY, we benchmark it in an obstacle-rich environment (Fig. 7) against EGO-Swarm [24], EGO-Swarm2 [8], and SUPER [14]. Static cylindrical obstacles are placed randomly with radii sampled in $1.0\text{\,}\mathrm{m}$,$1.5\text{\,}\mathrm{m}$\$\,] and height $6.0\text{\,}\mathrm{m}$, occupying a $300\text{\,}\mathrm{m}$\times$40\text{\,}\mathrm{m}$$ area. The agent starts at $(0,0,3)\,$\mathrm{m}$\$ and the goal is $(305,0,3)\,$\mathrm{m}$\$. We constrain velocity, acceleration, and jerk to $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, $a_{\max}=$10.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, and $j_{\max}=$30.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{3}$\$. For global planning we run A*, then build an SFC [3] for local optimization. The intermediate waypoints for A* are projected from the terminal goal onto the local occupancy map centered on the drone. The average computation time for A* and SFC generation is $0.26\text{\,}\mathrm{m}\mathrm{s}$ and $0.48\text{\,}\mathrm{m}\mathrm{s}$, respectively.
For soft constraints we use a smooth hinge $\phi_{\mu}(\cdot)$. Time integrals are evaluated by trapezoidal quadrature with $\kappa$ samples per segment unless a closed form is available. SFCs are modeled per segment $s$ as the intersection of halfspaces
$\text{SFC}_{s}\;\triangleq\;\big\{\mathbf{x}\in\mathbb{R}^{3}\;\big|\;\mathbf{n}_{s,h}^{\top}\mathbf{x}\leq b_{s,h}-C_{\text{SFC}},\ \ h=1,\dots,H_{s}\big\},$
@@
where $J_{\text{SFC}}\;=\;\sum_{s}\int_{0}^{T_{s}}\sum_{h=1}^{H_{s}}\phi_{\mu}\!\big(\,\mathbf{n}_{s,h}^{\top}\mathbf{x}(t)-b_{s,h}+C_{\text{SFC}}\,\big)\,dt,$ penalizes leaving the SFC, $J_{v}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{v}(t)\|^{2}-v_{\max}^{2})\,dt$, $J_{a}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{a}(t)\|^{2}-a_{\max}^{2})\,dt$, and $J_{j}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{j}(t)\|^{2}-j_{\max}^{2})\,dt$. Note that the jerk smoothness cost $\tilde{J}_{\text{smooth}}$ uses the closed form given in Eq. (6), which avoids per-sample recomputation. MIGHTY’s weights are $w_{T}=$5\text{\times}{10}^{2}$\$, $w_{\mathrm{smooth}}=${10}^{-1}$\$, $w_{\text{SFC}}=${10}^{3}$\$, $w_{v}=${10}^{3}$\$, $w_{a}=${10}^{3}$\$, $w_{j}=${10}^{3}$\$. We run 10 trials per method. Metrics: $R_{\mathrm{succ}}$ [%] (success rate; collision-free, reaches goal); $T_{\mathrm{opt}}$ [ms] (local optimization time); $T_{\mathrm{total}}$ [ms] (total planning time); $T_{\mathrm{trav}}$ [s] (travel time); $L_{\mathrm{path}}$ [m] (path length); $S_{\mathrm{jerk}}=\!\int\!\lVert\mathbf{j}(t)\rVert\,dt$ [m/s2] (smoothness); $\bar{S}_{\mathrm{jerk}}=\sqrt{\tfrac{1}{T}\int\!\lVert\mathbf{j}(t)\rVert^{2}\,dt}$ [m/s3] (RMS jerk; time-normalized); $\rho_{\mathrm{vel}},\,\rho_{\mathrm{acc}},\,\rho_{\mathrm{jerk}}$ [%] (velocity/acceleration/jerk violations). We report $\bar{S}_{\mathrm{jerk}}$ to account for differing travel times across methods.